OpenStudy (crashonce):

The sum of an infinite GP is 15. The sum of the squares of the terms is 45. Find the first term @iambatman

OpenStudy (crashonce):

no solution probided

OpenStudy (dan815):

well write it out

OpenStudy (dan815):

u have 2 eqn and 2 unknowns

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} ax^n = \frac{ a }{ 1-x }\] set this equal to 15 find the pattern set the second equation = 45 and go from there.

OpenStudy (anonymous):

I suppose you'd start with:\[ 15 = \sum_{k=1}^{\infty}ar^{k-1}=\frac{a}{1-r} \]

OpenStudy (crashonce):

what about the squared part not sure about that

OpenStudy (anonymous):

\[\frac{ a^2 }{ 1-x^2 } = 45\] do you mean this?

OpenStudy (anonymous):

If you squared the terms, you'd get \[ 45 =\sum_{k=0}^{\infty }( ar^{k-1} )^2 =\sum_{k=0}^{\infty }a^2(r^2)^{k-1} = \frac{a^2}{1-r^2} \]

OpenStudy (anonymous):

I'll let wio handle this :D

OpenStudy (dan815):

xD it was solved at 2 eqn and 2 unknown

OpenStudy (anonymous):

\[ 45 = \frac{a}{1+r}\cdot \frac{a}{1-r} = \frac{a}{1+r}\cdot 15 \]

OpenStudy (crashonce):

oooooooo right

OpenStudy (anonymous):

Hahaha.

OpenStudy (crashonce):

can u solve the rest for me @wio

OpenStudy (anonymous):

I probably could.

OpenStudy (anonymous):

First think you want to do is isolate the \(r\) so you can sub it into the other formula.

OpenStudy (crashonce):

solve it please

OpenStudy (anonymous):

Why don't you finish it off, what is r?

OpenStudy (anonymous):

I meant to use r instead of x..

OpenStudy (crashonce):

(15-a)/15

OpenStudy (anonymous):

I don't know what that means

OpenStudy (crashonce):

\[\frac{ 15-a }{ 15 }\]

OpenStudy (anonymous):

I'm not sure what you mean, I got r = 2/3?

OpenStudy (crashonce):

dw i meant at first, not when solved

OpenStudy (anonymous):

Well you have r now lol, and a should be obvious so you can get the first term now :).

OpenStudy (crashonce):

5 lol

OpenStudy (anonymous):

Yeah :D